# How does Architecture Influence the Base Capabilities of Pre-trained Language Models? A Case Study Based on FFN-Wider Transformer Models

### v1

@@ 7 From FFN-Wider Transformers to MoE Architecture Transformers
The MoE Transformer is a practical architecture that uses a mix of experts to replace the FFN layer. It allows different representations to be processed by different experts, which enables expanding the capacity with lower computation expense.
@@
However, base capability decline also be observed in MoE Transformers.  (Fedus et al., 2022) found MoE models perform lower on SuperGLUE compared to vanilla models when achieving same pre-training level. Similar issue can also be analyzed from the results of (Artetxe et al., 2022).
@@
We found the MoE layer can be seen as an enhanced version of the FFN layer, and the MoE Transformers share similarities with the FFN-Wider Transformers we previously studied. Therefore, we believe the previous explanations could also apply to MoE Transformers. Consequently, we directly transplant our CEA to MoE Transformers, resulting in a new MoE architecture.
We conducted experiments on BERT. The number of experts was set to 64, and each representation is processed by one expert. We pre-trained vanilla BERT, vanilla MoE and new architecture MoE models, aligning their pre-training performances. Representative results are shown in Table 3.
The results indicate the base capabilities of the new architecture MoE have indeed improved to some extent. However, it did not achieve comprehensive improvements. This suggests the reasons for base capability decline might be multifaceted. There may be other issues within MoE leading to this decline and our explanation only covers one aspect, which warrants further investigation. But overall, this still indicates our analysis can offer valuable guidance for architecture improvement.
@@ 8 Conclusion
9 Limitations
Although this work provides some insights into how architecture influences the base capabilities of models and conducts many targeted experiments, there are still some limitations.
1. This work mainly analyzes and experiments with models that use language modeling as the pre-training objective, lacking experiments on models with other pre-training objectives. Hence, the conclusions are limited to pre-trained language models. We focus mainly on the vanilla Transformer and specific variants, without experimenting on other complex variants, limiting the conclusions to the models tested. Therefore, the current applicability of our findings is relatively narrow, and we consider conducting more experiments in the future.
2. Although this work involves experiments on models of two different scales and conducts rigorous verification, our resources limit us from pre-training and experimenting with larger-scale models. This indirectly prevents us from conducting complex few-shot and more crucial instruction tuning experiments. We plan to improve these aspects through collaboration in the future.
3. Although preliminary experiments are conducted on the MoE Transformers, the results obtained are still very limited, and we consider conducting more in-depth research subsequently.
@@ Appendix A Other Evaluation Schemes
Parameter numbers alignment   It imposes restrictions on the model that make it inapplicable to any two different architecture models. More critically, this scheme has clear counterexamples, as it is not suitable for ALBERT models Lan et al. (2020). Under the same parameter numbers, ALBERT models exhibit strong base capabilities, but this is due to the large computational load, not much related to architecture inductive biases.
@@
Computational load alignment   It imposes restrictions on the model that make it inapplicable to any two different architecture models. More critically, this scheme also has clear counterexamples, as it is not suitable for MoE models Lepikhin et al. (2021); Fedus et al. (2022). Under the same computational load, MoE models demonstrate strong base capabilities, but this is due to large parameter numbers and model capacity, not much related to architecture inductive biases.
@@
Although these schemes have shortcomings, they also have reasonable aspects. To indisputably prove base capability improvements come from our architecture changes, we also conducted experiments that align pre-training performance, pre-training steps, parameter numbers and computational load simultaneously, as detailed in Appendix H.
@@
Appendix B Model Specifications and Pre-training Procedures
@@
All experiments were conducted in English only. We concurrently pre-trained both BERT and GPT models, where the architecture design of the BERT model adheres to the work of  Devlin et al. (2019), and the GPT model structure follows that established by  Radford et al. (2018). Both models employ a post layer normalization scheme.
@@
We trained small and large scales of both BERT and GPT. The small-scale models have a hidden dimension of 128, 12 layers, and 2 attention heads; the large-scale models possess a hidden dimension of 768, 12 layers, and 12 attention heads. In vanilla BERT and GPT, the FFN intermediate dimension is 4 times the hidden dimension, while the ratio of FFN-Wider models is 32 times. The small-scale vanilla models have 6.3M Parameters and the small-scale other models have 17.3M Parameters. The large-scale vanilla models have 110M Parameters and the large-scale other models have 506M Parameters. All models have a maximum sequence length of 128 and utilize the BERT vocabulary released by  Devlin et al. (2019), comprising 30,522 tokens. We used PyTorch [^1] https://pytorch.org/ and transformers [^2] https://github.com/huggingface/transformers libraries.
@@
The BERT model employs a masked language modeling task for pre-training, masking 15% of tokens in the sequence—80% are replaced with [MASK], 10% with random tokens, and 10% remain unchanged. Differing from  Devlin et al. (2019), we removed the next sentence prediction task, using long and continuous text for pre-training inputs and applying different masking schemes to the same input sequence in various epochs. The GPT model is pre-trained using a language modeling task without additional special settings.
@@
The maximum epoch set for all model pre-training is 40, but in practice, it was not reached; mid-training checkpoints were used for alignment and experimentation. All models used the Adam optimizer Kingma and Ba (2015) for pre-training, with a learning rate of 1e-4, $\beta_{1}=0.9$, $\beta_{2}=0.999$, L2 weight of 0.01, a warm-up over the first 10,000 steps, followed by linear decay. The small-scale models were pre-trained with a batch size of 512 on four Nvidia Tesla V100s, and total GPU days are approximately 55 days; the large-scale models with a batch size of 1024 on four Nvidia Tesla A100s, and total GPU days are approximately 87 days.
@@
To fairly compare the base capabilities of different architecture models, all models were pre-trained from scratch. However, as mentioned earlier, due to limited computational resources, our pre-training steps generally fell short of those in the original papers Devlin et al. (2019); Radford et al. (2018, 2019); Brown et al. (2020), which may result in discrepancies in downstream task performance compared to the original research.
Appendix C Fine-tuning Settings
@@
The maximum number of training epochs during fine-tuning was 10, with a batch size of 32. The optimizer was Adam Kingma and Ba (2015), with a warmup ratio of 0.06, a linearly decaying learning rate, and a weight decay of 0.01. We reported the average performance of multiple runs.
For the small-scale (H=128) models, we observed models with a wider FFN might overfit when fully fine-tuned with all parameters. Thus, we performed both full parameter fine-tuning and efficient fine-tuning based on adapters Houlsby et al. (2019) for all models, choosing the better result of the two for reporting. For full parameter fine-tuning, the learning rates were {1e-5, 2e-5, 5e-5}; for efficient fine-tuning, the adapter size was 128, with learning rates of {1e-4, 2e-4, 3e-4}. For the large-scale (H=768) models, we directly conducted full parameter fine-tuning for all models, with learning rates of {1e-5, 2e-5, 5e-5}.
Appendix D Few-shot Learning Settings
In our study on GPT models, we conducted few-shot learning experiments on multiple datasets. Since small-scale (H=128) models have limited capabilities and struggle with few-shot learning, we focused our experiments only on large-scale (H=768) models for 0-shot and 1-shot learning.
We followed the method of  Brown et al. (2020), which involves transforming the classification into comparisons of probability magnitudes.
Specifically, the unified format of these datasets involves selecting one option from multiple choices, given a context. We concatenated the context with different options and compared their respective generation probabilities. This comparison could be based on either the probability of the option alone or the entire sequence text, with a choice of normalizing the probabilities by length or not.  Brown et al. (2020) mentioned a method of normalizing the probability of each option unconditionally, which we also included in our options. We then experimented with all possible combinations of these choices for each dataset, determined the best combination for each, and reported the results of all models under these optimal conditions.
For 1-shot experiment, we randomly selected one demonstration each time, repeated the experiment 10 times, and reported the average results.
Appendix E Mutual Information Estimate
For estimating the mutual information between intermediate representations and target tokens, we adopted the method proposed by  (Voita et al., 2019), which primarily involves transforming the representations into discrete variables through clustering and then calculating mutual information.
@@ Appendix E Mutual Information Estimate
Appendix F Additional Transformation Function Contribution in BERT
@@
Appendix G Detailed Results (Main Setting: Pre-training Performance Alignment)
In Section 6.2, we present the brief experimental results in Table 1 and 2, and here we provide the corresponding detailed results.
@@ Appendix G Detailed Results (Main Setting: Pre-training Performance Alignment)
Appendix H Detailed Results (Extra Setting: Pre-training Performance Alignment & Pre-training Steps Alignment)
The brief results are shown in Table 4 and 5, and the corresponding detailed results are as follows:
Table 4: The results of BERT models (Pre-training Steps Alignment & Pre-training Performance Alignment).
@@ Appendix H Detailed Results (Extra Setting: Pre-training Performance Alignment & Pre-training Steps Alignment)
Table 5: The results of GPT models (Pre-training Steps Alignment & Pre-training Performance Alignment).

### v3

@@ 7 From FFN-Wider Transformers to MoE Transformers
The Mixture of Experts (MoE) Transformer [18, 8] is a practical architecture that introduces a mix of experts, which enables expanding the capacity with lower computation expense.
@@
However, base capability decline also be observed in MoE Transformers.  [8] found MoE models perform lower on SuperGLUE compared to vanilla models when achieving same pre-training level. Similar issue can also be analyzed from the results of [1].
@@
We found the MoE layer can be seen as an enhanced version of the FFN layer. Therefore, we believe the previous explanations could also apply to MoE Transformers. Consequently, we directly transplant our CEA to MoE Transformers, resulting in a new MoE architecture.
We selected a 1.3B GPT model as the backbone model, which incorporates Rotary Embedding [30] and RMSNorm [37] on the GPT3 [3]. Then, we added a MoE layer with 64 experts (top-1 activation) before the FFN layer, resulting in a 14B parameter MoE baseline model. Finally, our improved version with CEA involves transforming the MoE layer into an Inner-MoE layer within the MHA, while retaining the original FFN layer as the Outer-FFN layer. To accommodate FlashAttention-2 [6], we simplified the direct pathway. Instead of preventing transformation leakage by replacing the key and value at current position, we chose to directly mask the current position to achieve the same effect. The specifications, pre-training procedures and few-shot learning procedures are detailed in Appendix C and E.
We conducted pre-training on CC and C4 subsets within SlimPajama [29], training all models from scratch with 100B tokens. Due to the de-duplication across subsets achieved by SlimPajama, we used the remaining subsets as out-of-distribution sets. We carried out comprehensive out-of-distribution language modeling and few-shot learning tests, with the results presented in Table 1.
From the results, it is evident that our method significantly enhances the base capabilities of the MoE model. This fully demonstrates the effectiveness of our analysis and improvement methods.
8 Limitations
This work mainly analyzes the models that use language modeling as the pre-training objective, lacking experiments on models with other pre-training objectives. Hence, the conclusions are limited to pre-trained language models. Therefore, the current applicability of our findings is relatively narrow, and we consider conducting more experiments in the future.
@@ 9 Conclusion
Acknowledgments
This work was supported by the Fundamental Research Funds for the Central Universities (project number: 2022FRFK060002), the National Key RD Program of China via grant 2021YFF0901602 and the National Natural Science Foundation of China (NSFC) via grant 62176078.
@@ Appendix A Other Evaluation Schemes
Parameter numbers alignment   It imposes restrictions on the model that make it inapplicable to any two different architecture models. More critically, this scheme has clear counterexamples, as it is not suitable for ALBERT models [17]. Under the same parameter numbers, ALBERT models exhibit strong base capabilities, but this is due to the large computational load, not much related to architecture inductive biases.
@@
Computational load alignment   It imposes restrictions on the model that make it inapplicable to any two different architecture models. More critically, this scheme also has clear counterexamples, as it is not suitable for MoE models [18, 8]. Under the same computational load, MoE models demonstrate strong base capabilities, but this is due to large parameter numbers and model capacity, not much related to architecture inductive biases.
@@
Although these schemes have shortcomings, they also have reasonable aspects. To indisputably prove base capability improvements come from our architecture changes, we also conducted experiments that align pre-training performance, pre-training steps, parameter numbers and computational load simultaneously, as detailed in Appendix I.
In addition, the experiment with MoE transformers also adopts a stricter setting similar to the above. In the comparative experiments between the vanilla MoE model and the improved MoE model, not only is pre-training performance aligned, but pre-training steps, parameter numbers, and computational load are also kept consistent.
Appendix B Evaluation Tasks
B.1 FFN-Wider Transformers
All experiments with FFN-Wider transformers follow the unified settings described here:
The pre-training corpus is consistent with that of BERT [7], namely Wikipedia and BooksCorpus. We partition a portion of the corpus to serve as an in-distribution development set, and use it to align pre-training performance.
@@ B.1 FFN-Wider Transformers
We conduct experiments on four specifications: BERT (H=128), BERT (H=768), GPT (H=128) and GPT (H=768). The specifications and pre-training procedures are detailed in Appendix C.
For out-of-distribution language modeling capability, we evaluate on the development set of Pile dataset [9].
For transfer learning capability, we only evaluate BERT models on GLUE [34], SuperGLUE [33], HellaSwag [36], PIQA [2], OpenBookQA [20], ARC Easy & Challenge [5] and WinoGrande [27]. The experimental settings are detailed in Appendix D.
For few-shot learning capability, we only evaluate GPT models. Limited by the maximum sequence length (128) of our pre-trained models, we conduct 0-shot and 1-shot experiments on HellaSwag, PIQA, OpenBookQA, ARC Easy & Challenge, WinoGrande, Winograd [19], COPA [26] and StoryCloze [21]. The experimental settings are detailed in Appendix E.
B.2 MoE Transformers
All experiments with MoE transformers follow the unified settings described here:
The pre-training corpus is the CC and C4 subsets within the SlimPajama dataset [29]. The specifications and pre-training procedures are detailed in Appendix C.
For out-of-distribution language modeling capability, due to the de-duplication across subsets achieved by SlimPajama [29], we used the remaining subsets (Arxiv, Book, Github, Stack and Wiki) as out-of-distribution test sets.
For few-shot learning capability, we conduct 0-shot experiments on LAMBADA [23], and conduct 5-shot experiments on MMLU [10], OpenBookQA, ARC Easy & Challenge, BoolQ [4], RACE Middle & High [16], SIQA [28], SCIQ [35], HellaSwag, COPA, PIQA, StoryCloze, WinoGrande and Winograd. The experimental settings are detailed in Appendix E.
Appendix C Model Specifications and Pre-training Procedures
C.1 FFN-Wider Transformers
@@
All experiments were conducted in English only. We concurrently pre-trained both BERT and GPT models, where the architecture design of the BERT model adheres to the work of  [7], and the GPT model structure follows that established by  [24]. Both models employ a post layer normalization scheme.
@@
We trained small and large scales of both BERT and GPT. The small-scale models have a hidden dimension of 128, 12 layers, and 2 attention heads; the large-scale models possess a hidden dimension of 768, 12 layers, and 12 attention heads. In vanilla BERT and GPT, the FFN intermediate dimension is 4 times the hidden dimension, while the ratio of FFN-Wider models is 32 times. The small-scale vanilla models have 6.3M Parameters and the small-scale other models have 17.3M Parameters. The large-scale vanilla models have 110M Parameters and the large-scale other models have 506M Parameters. All models have a maximum sequence length of 128 and utilize the BERT vocabulary released by  [7], comprising 30,522 tokens. We used PyTorch [^1] https://pytorch.org/ and transformers [^2] https://github.com/huggingface/transformers libraries.
@@
The BERT model employs a masked language modeling task for pre-training, masking 15% of tokens in the sequence—80% are replaced with [MASK], 10% with random tokens, and 10% remain unchanged. Differing from  [7], we removed the next sentence prediction task, using long and continuous text for pre-training inputs and applying different masking schemes to the same input sequence in various epochs. The GPT model is pre-trained using a language modeling task without additional special settings.
@@
The maximum epoch set for all model pre-training is 40, but in practice, it was not reached; mid-training checkpoints were used for alignment and experimentation. All models used the Adam optimizer [15] for pre-training, with a learning rate of 1e-4, $\beta_{1}=0.9$, $\beta_{2}=0.999$, L2 weight of 0.01, a warm-up over the first 10,000 steps, followed by linear decay. The small-scale models were pre-trained with a batch size of 512 on four Nvidia Tesla V100s, and total GPU days are approximately 55 days; the large-scale models with a batch size of 1024 on four Nvidia Tesla A100s, and total GPU days are approximately 87 days.
@@
To fairly compare the base capabilities of different architecture models, all models were pre-trained from scratch. However, as mentioned earlier, due to limited computational resources, our pre-training steps generally fell short of those in the original papers [7, 24, 25, 3], which may result in discrepancies in downstream task performance compared to the original research.
C.2 MoE Transformers
All experiments were conducted in English only.
We selected the GPT model as the backbone model (Vanilla GPT 1.3B), which incorporates Rotary Embedding [30] and RMSNorm [37] on the GPT3 [3]. Then, we added a MoE layer before the FFN layer, resulting in our MoE baseline model (Vanilla MoE 14B). Finally, our improved version with CEA (MoE 14B w/ CEA) involves transforming the MoE layer into an Inner-MoE layer within the MHA, while retaining the original FFN layer as the Outer-FFN layer. To accommodate FlashAttention-2 [6], we simplified the direct pathway. Instead of preventing transformation leakage by replacing the key and value at current position, we chose to directly mask the current position to achieve the same effect.
The Vanilla GPT 1.3B has a hidden dimension of 4,096, 6 layers and 32 attention heads; the Vanilla MoE 14B has a hidden dimension of 4,096, 6 layers, 32 attention heads and 6 MoE layers with 64 experts (top-1 activation, intermediate dimension is 4,096) before the corresponding FFN layers; the MoE 14B w/ CEA is similar to the Vanilla MoE 14B, just the position of the parameters has changed. All models have a maximum sequence length of 2,048 and utilize the Mixtral vocabulary released by  [13], comprising 32,000 tokens. We used PyTorch, transformers and fastmoe [^3] https://github.com/laekov/fastmoe libraries.
These models are pre-trained using a language modeling task without additional special settings. We conducted pre-training on CC and C4 subsets within SlimPajama [29], training all models from scratch with 100B tokens. Due to the de-duplication across subsets achieved by SlimPajama, we used the remaining subsets as out-of-distribution sets. All models used the Adam optimizer [15] for pre-training, with a learning rate of 5e-4, $\beta_{1}=0.9$, $\beta_{2}=0.95$, L2 weight of 0.01, a warm-up over the first 2,500 steps, followed by linear decay. All models were pre-trained with a batch size of 256 on 8 Nvidia Tesla A100 80G cards, and total GPU days are approximately 280 days.
Appendix D Fine-tuning Settings
@@
The maximum number of training epochs during fine-tuning was 10, with a batch size of 32. The optimizer was Adam [15], with a warmup ratio of 0.06, a linearly decaying learning rate, and a weight decay of 0.01. We reported the average performance of multiple runs.
For the small-scale (H=128) models, we observed models with a wider FFN might overfit when fully fine-tuned with all parameters. Thus, we performed both full parameter fine-tuning and efficient fine-tuning based on adapters [12] for all models, choosing the better result of the two for reporting. For full parameter fine-tuning, the learning rates were {1e-5, 2e-5, 5e-5}; for efficient fine-tuning, the adapter size was 128, with learning rates of {1e-4, 2e-4, 3e-4}. For the large-scale (H=768) models, we directly conducted full parameter fine-tuning for all models, with learning rates of {1e-5, 2e-5, 5e-5}.
Appendix E Few-shot Learning Settings
In our study on GPT and MoE models, we conducted few-shot learning experiments on multiple datasets. Since small-scale (H=128) models have limited capabilities and struggle with few-shot learning, we focused our experiments only on large-scale (H=768) models for 0-shot and 1-shot learning and MoE models for 0-shot and 5-shot learning.
We followed the method of  [3], which involves transforming the classification into comparisons of probability magnitudes.
Specifically, the unified format of these datasets involves selecting one option from multiple choices, given a context. We concatenated the context with different options and compared their respective generation probabilities. This comparison could be based on either the probability of the option alone or the entire sequence text, with a choice of normalizing the probabilities by length or not.  [3] mentioned a method of normalizing the probability of each option unconditionally, which we also included in our options. We then experimented with all possible combinations of these choices for each dataset, determined the best combination for each, and reported the results of all models under these optimal conditions.
For 1-shot experiment, we randomly selected one demonstration each time, repeated the experiment 10 times, and reported the average results. For 5-shot experiment, we randomly selected five demonstrations each time (for MMLU, randomly shuffling the order of demonstrations), repeated the experiment 5 times, and reported the average results.
Appendix F Mutual Information Estimate
The formal definition of mutual information for two discrete random variables is as follows:
$\mathrm{I}(X;Y)=\sum_{y\in\mathcal{Y}}\sum_{x\in\mathcal{X}}P_{(X,Y)}(x,y)\log\biggl{(}\frac{P_{(X,Y)}(x,y)}{P_{X}(x)P_{Y}(y)}\biggr{)}$ (1)
where $X$ and $Y$ are discrete random variables, $P_{(X,Y)}$ is the joint probability mass function of $X$ and $Y$, and $P_{X}$ and $P_{Y}$ are the marginal probability mass functions of $X$ and $Y$ respectively.
For estimating the mutual information between intermediate representations and target tokens, we adopted the method proposed by  [32], which primarily involves transforming the representations into discrete variables through clustering and then calculating mutual information.
@@ Appendix F Mutual Information Estimate
Appendix G Additional Transformation Function Contribution in BERT
@@
Appendix H Detailed Results (Main Setting: Pre-training Performance Alignment)
In Section 6.2, we present the brief experimental results in Table 8 and 8, and here we provide the corresponding detailed results.
@@ Appendix H Detailed Results (Main Setting: Pre-training Performance Alignment)
Table 2: The results of BERT (Pre-training Steps Alignment & Pre-training Performance Alignment).
@@
Table 3: The results of GPT (Pre-training Steps Alignment & Pre-training Performance Alignment).
@@
Appendix I Detailed Results (Extra Setting: Pre-training Performance Alignment & Pre-training Steps Alignment)
The brief results are shown in Table 2 and 3, and the corresponding detailed results are as follows: